# Evaluating Dialect Robustness of Language Models via Conversation Understanding

### v2

@@ Appendix A Dataset Construction
Table 5 describes the example conversations from extended MD3 and their corresponding masked versions from M-MD3. We mask the turn where the guesser utters the target word to help with formulating our downstream tasks. We mask the target word by finding the exact match in the conversation as shown in the conversations from Table 5. In case of conversations where an exact match is not found (such as planets), we find the utterance that is most similar to the target word using the similarity score101010Described in the Metrics section of the main paper.. The rest of the conversation is then pruned to make the masked target word (represented by ‘[MASK]’) the last token in the conversation.
@@ Appendix B Transformation Issues
We transform conversations from en-IN into en-TR, by prompting111111The exact prompt can be found in the Methodology section of the main paper. gpt-4 to remove exaggerations and dialectal information. As mentioned in Table 6, a ‘typical’ transformed conversation maintains the semantic meaning but only differs from the original conversation grammatically. A ‘bad’ example deviates largely from the expected output, and such examples are excluded from the final set of conversations used in our evaluation.
@@ Appendix C Errors
Table 7 describes additional examples for all identified error types121212defined in the Error Analysis section of the main paper.. As mentioned, each conversation can be classified under multiple error types. For example, the conversation about the target word–‘Ryan Reynolds’ is classified as CC, but can also be classified as PF.
Type Target Word Conversation Model Prediction AD Fisherman Describer: Okay. Okay. A. guy um wants to um okay. Guy catching something in the water. Fish Guesser: [MASK] WD Atlantic Ocean Describer: One of the. of the of world. Kanyakumari Guesser: Of the seventh wonder of the world. Taj mahal? Is it regarding sea? Describer: No no no the. Towards the bottom of India. Guesser: Is it regarding Describer: what we have? Guesser: [MASK] BDD Russian Language Describer: Ok. Ah Largest continent in the world Russian Guesser: Ok. Describer: Ah like area wise. Which country? Guesser: Largest. vast area. vast area? Russia but. Describer: We need to add N over there at the end. Guesser: Russian Describer: We speak Guesser: What they speak? Describer: Yeah. Ok. Guesser: [MASK] CC Ryan Reynolds Describer: It is like. One of the. Pen name. which we used in school school days. Flair Guesser: Cello point pen. Fine Grip Describer: No no no Guesser: Reynolds Describer: Uh yeah yeah Guesser: This is a second word or first word. Describer: Yeah this is second word Guesser: First word is. Name Describer:Yeah name related to the same Guesser: [MASK] PF Steve Jobs Describer: Ok. He is a famous person and he is a. a. for. what we call? . Um now it is a. Its. giving competition to Android. what we call? Steve Guesser: ok. so he is the fond ok sorry Describer: he is a founder of so and so company. Its a U. S. company Guesser: so it is giving competition to Android means Google ok.. So Describer: and he is the founder of that company Guesser: [MASK] ERR Podium Describer: Okay um. uh. well I isn’t sure I’m not sure but uh letting are seeing. Well it’s like preacher are churching. I am standing behind this. uh. in in used for speaker. Pulpit Guesser: [MASK]

### v3

@@ Appendix A Dataset Construction
Table 5 describes the example conversations from extended MD3 and their corresponding masked versions from M-MD3. We mask the turn where the guesser utters the target word to help with formulating our downstream tasks. We mask the target word by finding the exact match in the conversation as shown in the conversations from Table 5. In case of conversations where an exact match is not found (such as planets), we find the utterance that is most similar to the target word using the similarity score111111Described in Section 3.2 of the main paper.. The rest of the conversation is then pruned to make the masked target word (represented by ‘[MASK]’) the last token in the conversation.
@@ Appendix B Transformation Issues
We present examples of transformation issues faced while creating en-TR in Table 6. We create en-TR by prompting121212The exact prompt can be found in Section 2 of the main paper. gpt-4 to remove exaggerations and dialectal information from en-IN conversations. Table 7 presents examples of similar issues faced while creating en-MV using Multi-VALUE. As mentioned in Tables 6 and 7, a ‘typical’ transformed conversation maintains the semantic meaning but only differs from the original conversation grammatically. A ‘bad’ example deviates largely from the expected output. An ‘erroneous’ example is a result of Multi-VALUE not being able to transform a conversation from en-US. Both ‘bad’ and ‘erroneous’ examples are excluded from the final set of conversations used in our evaluation.
@@
Type en-US en-MV Typical Describer: Perfect. Oh! (We) earn this. We go to our jobs. Describer: Perfect. Oh! () [are] earn[ing] this. We [are] go[ing] to our jobs. Guesser: Money Guesser: Money Error Describer: This person. is in. oh films. It’s a man. He’s um. famous for a fine show in the ’80s. None Guesser: Um. what else is he in? Describer: He’s in a lot of things. I think he’s kind of short. Some people think that he looks nice but I don’t. Mhm. Mish Mission Impossible. Mission Impossible. Guesser: Tom Cruise
@@ Appendix C Errors
Table 8 describes additional examples for all identified error types131313Defined in Section 4.2 of the main paper.. As mentioned, each conversation can be classified under multiple error types. For example, the conversation about the target word–‘Ryan Reynolds’ is classified as CC, but can also be classified as PF.
Type Target Word Conversation Model Prediction AD Fisherman Describer: Okay. Okay. A. guy um wants to um okay. Guy catching something in the water. Fish Guesser: [MASK] Mike Tyson Describer: Big guy that punched people out and he had a little bit of a lisp. Darth Wader Guesser: [MASK] WD Atlantic Ocean Describer: One of the. of the of world. Kanyakumari Guesser: Of the seventh wonder of the world. Taj mahal? Is it regarding sea? Describer: No no no the. Towards the bottom of India. Guesser: Is it regarding Describer: what we have? Guesser: [MASK] Beg Describer: Um so if you don’t have any money uh you may stand on the corner. Panhandle Guesser: [MASK] BDD Russian Language Describer: Ok. Ah Largest continent in the world Russian Guesser: Ok. Describer: Ah like area wise. Which country? Guesser: Largest. vast area. vast area? Russia but. Describer: We need to add N over there at the end. Guesser: Russian Describer: We speak Guesser: What they speak? Describer: Yeah. Ok. Guesser: [MASK] Cold War Describer: This is a two-word term. The first word is a common illness that causes a runny nose. War Guesser: Cold. Describer: Yes that’s the first word. The second word refers to a conflict between two countries. Guesser: [MASK] CC Ryan Reynolds Describer: It is like. One of the. Pen name. which we used in school school days. Flair Guesser: Cello point pen. Fine Grip Describer: No no no Guesser: Reynolds Describer: Uh yeah yeah Guesser: This is a second word or first word. Describer: Yeah this is second word Guesser: First word is. Name Describer:Yeah name related to the same Guesser: [MASK] Mark Wahlberg Describer: Okay. Um. He was the original. of the Funky Bunch. But then he stopped music. Marky Guesser: [MASK] PF Steve Jobs Describer: Ok. He is a famous person and he is a. a. for. what we call? Um now it is a. Its. giving competition to Android. what we call? Steve Guesser: ok. so he is the fond ok sorry Describer: he is a founder of so and so company. Its a U. S. company Guesser: so it is giving competition to Android means Google ok.. So Describer: and he is the founder of that company Guesser: [MASK] Kanye West Describer: All right. Clint Guesser: How do you wanna skip that one Describer: All right. Now um. This guy he um. He just bought a ranch in Wyoming. Guesser: [MASK] ERR Podium Describer: Okay um. uh. well I isn’t sure I’m not sure but uh letting are seeing. Well it’s like preacher are churching. I am standing behind this. uh. in in used for speaker. Pulpit Guesser: [MASK]